Just the answer please, no need to show the work.

Suppose there are two goods, the prices of both goods are positive and a consumer's income is also positive. If the consumer's income doubles and the price of both goods triple,

the consumer's budget line gets steeper and shifts inward.

the slope of the consumer's budget line does not change but the budget line shifts outward away from the origin.

the slope of the consumer's budget line does not change but the budget line shifts inward toward the origin.

the consumer's budget line gets steeper and shifts outward.
Answer #1

Answer

Budget Line is given by :

Px*X + Py*Y = M--------------(1) where Px = price of x, Py = Price of Y and M = Income

=> Y = (M - Px*X)/Py

Slope = dY/dX = -(Px/Py)

Now New Price is 3Px and 3Py and New Income = 2M

Thus New Budget Line is given by :

3Px*X + 3Py*Y = 2M

=> Px*X + Py*Y = (2/3)M ---------------------------------(2)

=> Y = (2M - 3Px*X)/(3Py)

=> Slope = dY/dX = -(3Px)/(3Py) = -(Px/Py)

Hence we can see that slope remains same.

Comparing (1) and (2) that we have decrease in Income(real) and hence Budget line will shift inwards and slope remains same.

Hence, the correct answer is (c) the slope of the consumer's budget line does not change but the budget line shifts inward toward the origin.